# DISCO: Diversifying Sample Condensation for Efficient Model Evaluation

### v1

@@ 5 Experiments
Evaluation. We evaluate DISCO and baseline approaches using two complementary metrics. First, the Mean Absolute Error (MAE) of the model accuracies, reported as percentage points (%p), captures the absolute error of accuracy prediction. Second, to assess the consistency of the relative ordering of models, we report the Spearman rank correlation (Rank) in model ranking between the true and estimated model performances.
@@ 5.1 Main results
Table 1: DISCO achieves state-of-the-art test-set compression by using model signatures combined with PDS for accurate performance prediction. Compression of MMLU, HellaSwag (HS), Winogrande (WG) and ARC datasets by DISCO (ours), tinyBenchmarks, Metabench, and other baselines. For each dataset, we reduce the test set to 100 data points (except for Metabench, see below), achieving inference cost reduction of 99.3% and 99.0%, on MMLU and HS, respectively. Sig. + RF/kNN stands for model signature with Random Forest/kNN prediction (§ 4.2.2). Mean absolute error (MAE) is the %p difference in accuracy, and Rank is the Spearman rank correlation between the true model ranking and the estimated model ranking.
†  Results for Metabench are not directly comparable, as it requires more examples to converge: 150 datapoints for MMLU and ARC (+50%), 450 for HS (+350%), and 200 for WG (+100%).
@@
We measure the efficacy of DISCO in two steps: adopt model-signature approach on top of uniform random sample selections first, and then consider sampling according to predictive diversity scoring (PDS). Even without PDS, on uniform random samples, model signatures are achieving 1.81%p MAE and .933 rank correlation with Random Forest (RF), reaching the state-of-the-art performance with simple and practical ingredients. When PDS is further considered for sample selection, to diversify the model outputs, we achieve 1.07%p MAE and .987 rank correlation, demonstrating a significant leap from the prior state of the art from tinyBenchmarks Polo et al. (2024) from ICML 2024.
Figure 4: True and estimated performance on MMLU. Scatter plot of performances of 40 models.
To provide an understanding of the distributional comparison of the true model performances and the estimated performances, we show a scatter plot in Figure 4. As signified by the high Pearson’s correlation coefficient at .986, the estimated performances closely follow the true performances.
Figure 5: MMLU performance estimation vs. compression rates. Mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation between the true model ranking and the estimated model ranking are shown. At 100 samples, the results are identical to Table 1. Main observations: DISCO hits a better efficiency-precision trade-off across all range of compression rates. For extreme compression rate, kNN is a better choice than random forest (RF).
Figure 5 shows the performance against varying degrees of the test set reduction. We observe that the ranking of estimated evaluation methodologies does not change much across a wide range of degrees of reduction. In particular, our DISCO is consistently the best method across all ranges of number of samples involved. For the extreme rates of compression, at 10 samples, the non-parametric performance predictor of kNN yields better performance than the parametric Random Forest, suggesting that non-parametric approaches may be more suitable at extreme compression.
5.2 Factors analysis
We analyse the impact of several design choices involved in our DISCO on MMLU dataset. See Table 2 for an overview.
@@ 5.2 Factors analysis
For this purpose, we introduced the chronological split in §5. We examine the impact of this model splitting on the result. We observe that our DISCO is robust to the choice of splitting strategy. Chronological splitting yields a rank correlation of .987, which is nearly identical to the .986 obtained with uniform splitting (Table 2 (a)).
Stratification. We measure the efficacy of the stratification strategy in (Polo et al., 2024) where equal numbers of anchor points are selected from each of 57 tasks in MMLU dataset (Table 2 (b)). We find that stratification (.978) is not effective when data points are sampled according to PDS (.987).
@@
Prediction model. We compare consider a wide range of prediction models (Table 2 (e)). Random Forest achieves the highest rank correlation of .987, outperforming all other methods.
@@ 5.3 Results for vision domain
Table 3: DISCO compression of ImageNet validation dataset. We evaluate the generalisation of our DISCO to the computer vision domain. The main metrics are mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation (Rank) between the true model ranking and the estimated model ranking. Main observations: (1) Same as for language experiments, model signature is an effective strategy for performance estimation. (2) Using PDS on top improves performance even more.
In this section, we give a quick overview of the DISCO applied to the vision domain. For detailed results, see § C.
Setup. We use ImageNet-1k (Russakovsky et al., 2015) with 1.28M images and 400 pretrained models from timm (Wightman, 2019), spanning convolutional (Krizhevsky et al., 2012) and transformer (Dosovitskiy et al., 2021) architectures (0.3M–300M parameters). Following the language domain, we adopt a chronological split with cutoff 5 April 2023 (88:12 train–test). Performance is evaluated using mean absolute error (MAE) and Spearman rank correlation.
Results. Our DISCO approach significantly compresses the ImageNet validation set by reducing it to just 100 data points, achieving an inference cost reduction of 99.8%. DISCO with uniform random sampling and random forest prediction on model signatures achieve 0.86%p MAE and .944 rank correlation, surpassing the baseline. Using a predictive diversity score (PDS) for data selection and a Random Forest for prediction, our method achieves a 0.63%p MAE and .969 rank correlation, substantially outperforming the baseline (Table 3). The results demonstrate that DISCO is effective in both language and vision domains.
@@ 6 Conclusion
We propose DISCO, which selects a small informative subset of the evaluation data that maximizes LLM disagreement and estimates model performance from predictions on it. DISCO cuts evaluation costs by over 99% with minimal error and consistently outperforms prior methods.
@@
Limitations. The main limitation of DISCO is robustness to distribution shifts in the model population. Shifts can arise from new architectures, training methods, or objectives, introducing patterns unseen during training and reducing estimator accuracy. Future work could address this with adaptive sample selection or periodic retraining on newer models.
@@ Acknowledgments
This work was supported by the Tübingen AI Center. AR thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for support. This research utilized compute resources at the Tübingen Machine Learning Cloud, DFG FKZ INST 37/1057-1 FUGG.
Appendix A Mutual Information and Jensen-Shannon Divergence

### v2

@@ 5.2 Setup
Metrics. We evaluate DISCO and baseline approaches using two complementary metrics. First, the Mean Absolute Error (MAE) of the model accuracies, reported as percentage points (%p), captures the absolute error of accuracy prediction. Second, to assess the consistency of the relative ordering of models, we report the Spearman rank correlation (Rank) in model ranking between the true and estimated model performances.
Models. Building on the TinyBenchmarks framework (Polo et al., 2024), we evaluate 424 large language models (LLMs) from Hugging Face’s Open LLM Leaderboard (Fourrier et al., 2024). The models cover GPT- (Radford et al., 2019), LLaMA- (Touvron et al., 2023), DeepSeek- (DeepSeek-AI et al., 2025), and BERT-style (Devlin et al., 2019) architectures, with model sizes ranging from 1.3 billion to 72 billion parameters.
Model split. DISCO is based on a meta-model approach where a predictor is constructed based on the model signatures of a pool of source models $\mathcal{F}$ and tested on a disjoint set of target models. This approach has traditionally been criticised for its dependency on the set of existing models: the predictor may fail to retain performance with unforeseen changes in future models. To address this concern, we introduce the chronological split, where the source models $\mathcal{F}$ consist of models published before January 13, 2024, and the meta test set consists of models after the cutoff date. The train-test ratio is 9:1.
@@ 5.3 Main results
Table 1: DISCO achieves state-of-the-art test-set compression by using model signatures combined with PDS for accurate performance prediction. Compression of MMLU, HellaSwag (HS), Winogrande (WG), and ARC datasets by DISCO (ours), tinyBenchmarks, Metabench, and other baselines. For each dataset, we reduce the test set to 100 data points (except for Metabench, see below), achieving inference cost reduction of 99.3% and 99.0%, on MMLU and HS, respectively. Sig. + RF/kNN stands for model signature with Random Forest/kNN prediction (§ 4.2.2). Mean absolute error (MAE) is the %p difference in accuracy, and Rank is the Spearman rank correlation between the true model ranking and the estimated model ranking.
†  Results for Metabench are not directly comparable, as it requires more examples to converge: 150 datapoints for MMLU and ARC (+50%), 450 for HS (+350%), and 200 for WG (+100%). Confidence intervals in App. D.
@@
Figure 4: True and estimated performance on MMLU. Scatter plot of the performances of 40 models.
We measure the efficacy of DISCO in two steps: adopt a model-signature approach on top of uniform random sample selections first, and then consider sampling according to predictive diversity scoring (PDS). Even without PDS, on uniform random samples, model signatures are achieving 1.81%p MAE and .933 rank correlation with Random Forest (RF), reaching the state-of-the-art performance with simple and practical ingredients. When PDS is further considered for sample selection, to diversify the model outputs, we achieve 1.07%p MAE and .987 rank correlation (see Appendix C for qualitative comparison of predicted ranks for DISCO vs direct evaluation), demonstrating a significant leap over the prior state of the art from tinyBenchmarks Polo et al. (2024) from ICML 2024.
To provide an understanding of the distributional comparison of the true model performances and the estimated performances, we show a scatter plot in Figure 4. As signified by the high Spearman’s correlation coefficient of .987, the estimated performances closely follow the true performances.
Figure 5: MMLU performance estimation vs. compression rates. Mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation between the true model ranking and the estimated model ranking are shown. At 100 samples, the results are identical to Table 1. Main observations: DISCO hits a better efficiency-precision trade-off across the entire range of compression rates. For an extreme compression rate, kNN is a better choice than random forest (RF).
Figure 5 shows the performance against varying degrees of the test set reduction. We observe that the ranking of estimated evaluation methodologies does not change much across a wide range of degrees of reduction. In particular, our DISCO is consistently the best method across all ranges of the number of samples involved. For the extreme rates of compression, at 10 samples, the non-parametric performance predictor of kNN yields better performance than the parametric Random Forest, suggesting that non-parametric approaches may be more suitable at extreme compression.
5.4 Factor analysis
We analyse the impact of several design choices involved in our DISCO on the MMLU dataset. See Table 2 for an overview.
@@ 5.4 Factor analysis
For this purpose, we introduced the chronological split in §5.2. We examine the impact of this model splitting on the result. We observe that our DISCO is robust to the choice of splitting strategy. Chronological splitting yields a rank correlation of .987, which is nearly identical to the .986 obtained with uniform splitting (Table 2 (a)).
Stratification. We measure the efficacy of the stratification strategy in (Polo et al., 2024), where equal numbers of anchor points are selected from each of 57 tasks in the MMLU dataset (Table 2 (b)). We find that stratification (.978) is not effective when data points are sampled according to PDS (.987).
@@
Prediction model. We consider a wide range of prediction models (Table 2 (e)). Random Forest achieves the highest rank correlation of .987, outperforming all other methods.
@@ 5.5 Results for vision domain
Lifelong Bench. Uniform Weighted 2.06 0.838
correctness sum
SSEPY Uniform Weighted 3.05 0.762
confidence sum
@@
Table 3: DISCO compression of ImageNet validation dataset. We evaluate the generalisation of our DISCO to the computer vision domain. We reduce the test set to 100 anchor points. The main metrics are mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation (Rank) between the true model ranking and the estimated model ranking. Main observations: (1) Same as for language experiments, model signature is an effective strategy for performance estimation. (2) Using PDS on top improves performance even more.
In this section, we give a quick overview of the DISCO applied to the vision domain. For detailed results, see Appendix J.
Setup. We use ImageNet-1k (Russakovsky et al., 2015) with 1.28M images and 400 pretrained models from timm (Wightman, 2019), spanning convolutional (Krizhevsky et al., 2012) and transformer (Dosovitskiy et al., 2021) architectures (0.3M–300M parameters). Following the language domain, we adopt a chronological split with a cutoff of 5 April 2023 (88:12 train–test). Performance is evaluated using mean absolute error (MAE) and Spearman’s rank correlation. The details on the baselines for the vision domain are in Appendix J.2.
Results. Our DISCO approach significantly compresses the ImageNet validation set by reducing it to just 100 data points, achieving an inference cost reduction of 99.8%. DISCO with uniform random sampling and random forest prediction on model signatures achieves 0.86%p MAE and .944 rank correlation, surpassing the baseline. Using a predictive diversity score (PDS) for data selection and a Random Forest for prediction, our method achieves a 0.63%p MAE and .969 rank correlation, substantially outperforming the baseline (Table 3). The results demonstrate that DISCO is effective in both language and vision domains.
DISCO (.969/0.63.969/0.63) outperforms Lifelong Bench. (Prabhu et al., 2024) (.838/2.06.838/2.06) and SSEPY (Fogliato et al., 2024) (.762/3.05.762/3.05) in both rank correlation and MAE. The conclusion from language experiments holds: instead of selecting anchor points with wide coverage of sample difficulty, one should focus on selecting the points on which models typically disagree.
@@ 6 Conclusion
We propose DISCO, which selects a small informative subset of the evaluation data and estimates model performance from predictions on it. DISCO cuts evaluation costs by over 99% with minimal error and consistently outperforms prior methods.
@@
Limitations. The main limitation of DISCO is robustness to distribution shifts in the model population. Shifts can arise from new architectures, training methods, or objectives, which introduce patterns unseen during training and reduce estimator accuracy. Future work could address this with adaptive sample selection or periodic retraining on newer models (see details in Appendix F).
We also discuss unsuitable tasks for DISCO. The main constraint is that DISCO requires predictive probabilities for several predefined answer choices for each question. These answer choices correspond to the classes in Proposition 1 in the original submission. That makes DISCO not suitable for open-ended generation tasks such as translation or summarisation. Applying DISCO to such tasks would first require defining sets of correct and incorrect outputs. We leave such experiments for future work.
@@ Acknowledgments
This work was supported by the Tübingen AI Center. AR thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for support. This research utilised compute resources at the Tübingen Machine Learning Cloud, DFG FKZ INST 37/1057-1 FUGG.
Appendix A Extended experimental setup
Datasets. We evaluate DISCO on four widely used language modelling benchmarks:
  • Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) question-answering dataset that covers 57 tasks about world knowledge and problem-solving ability.
  • HellaSwag (Zellers et al., 2019) dataset that focuses on commonsense natural language inference.
  • Winogrande (Sakaguchi et al., 2021): dataset of 273 expert-crafted pronoun resolution problems originally designed to be unsolvable for statistical models.
  • AI2 Reasoning Challenge (ARC) (Clark et al., 2018): question-answering dataset that contains only natural, grade-school science questions (authored for human tests) and requires knowledge and reasoning.
Appendix B Computational costs
We report the space–time complexity for the main stages of DISCO, as well as the cost of direct evaluation of a target model. The numbers correspond to a single H100 GPU and are extrapolated from evaluations of five diverse 32B LLMs on MMLU (standard deviation across 5 runs).
B.1 DISCO Pipeline Overview
The DISCO pipeline consists of two stages: an offline stage (run once) and an online stage (run for each new target model).
Offline Stage
  • Evaluate $M$ source models on the full test dataset ($M=385$ in this experiment)
  • Store source model outputs
  • Select 100 anchor points that maximise PDS/JSD
  • Concatenate outputs on anchor points to form model signatures
  • Train a predictor to estimate model performance on the full test dataset from these signatures
Online Stage
  • Evaluate one target model on the 100 anchor points
  • Store target model outputs
  • Concatenate to obtain the target model signature
  • Run the predictor to estimate performance on the full test dataset
For every target model, the anchor points and predictor trained offline are reused.
B.2 Cost Metrics
DISCO: offline stage $3284.05\pm 592.90$ GPU-hours
DISCO: online stage $0.07\pm 0.01$ GPU-hours
Direct evaluation $8.53\pm 1.54$ GPU-hours
Table 4: DISCO computation cost metrics (single H100 GPU, MMLU, 5 runs) compared to direct evaluation cost. Computation savings are computed as the difference between the direct evaluation cost and online computation cost, i.e., 8.530.07=8.468.53-0.07=8.46 GPU-hours (mean). This yields $8.46\pm 1.54$ GPU-hours saved per evaluated model.
Source outputs (offline) 86.54 MB
Source signatures (offline) 400 KB
Target outputs (online) 224.78 KB
Target signature (online) 1 KB
Table 5: DISCO storage requirements (offline stage for 400 source models and online stage for one target model).
The majority of the compute is required by the offline stage (3284 GPU-hours).
B.3 Break-Even Analysis: How Many Evaluations Justify DISCO Setup?
DISCO breaks even at 389 evaluations. Since each DISCO evaluation saves $8.46$ GPU-hours per model (vs. $8.53$ GPU-hours direct evaluation, minus $0.07$ GPU-hours online DISCO cost), the break-even point is:
$389=\frac{3284}{8.46}.$
In practice, hundreds of checkpoint evaluations naturally occur during model development. For example, a single OLMo-2-32B (OLMo et al., 2024) training run includes 753 checkpoints on Hugging Face, already exceeding break-even.
In some cases, there is no need to evaluate source models at all if offline predictions are downloaded from platforms such as open-llm-leaderboard.
B.4 Comparison to Alternative Approaches
We briefly remind the pipelines of the compared methods:
  • Selection: select a set of anchor points, i.e., a subset of the full test dataset based on different signals (random, IRT, model disagreement, etc.).
  • Prediction: estimate model performance on the full test dataset from outputs on anchor points.
That is why we use “Selection" and “Prediction" columns to explain the difference between methods. See § 4.1 for details on selection methods, and § 4.2 for details on prediction methods.
Method Selection Prediction Offline (GPU-h) Online (GPU-s)
Baseline – (use all samples) Direct eval $30739\pm 5514$
Baseline Random Direct eval $218\pm 39$
tinyBenchmarks Random gp-IRT $3284\pm 592$ $219\pm 39$
tinyBenchmarks Anchor-corr gp-IRT $3284\pm 592$ $219\pm 39$
tinyBenchmarks Anchor-IRT gp-IRT $3284\pm 592$ $219\pm 39$
DISCO High-PDS RF $3284\pm 592$ $218\pm 39$
DISCO High-PDS KNN $3284\pm 592$ $218\pm 39$
Table 6: Comparison to alternative approaches: offline (GPU-h) and online (GPU-s) costs, and computation savings (GPU-h). DISCO dominates all other methods.
The differences in online cost across methods are negligible (e.g., 219 vs. 218 GPU-seconds). Offline costs are equal up to rounding. Efficient evaluation methods allow for saving $\frac{(30739-218)}{30739}\cdot 100\%=99.3\%$ of evaluation cost in comparison to full evaluation.
Appendix C Qualitative Meaning of Rank Correlation Improvements
To justify the additional computation required for DISCO relative to direct evaluation, we illustrate how the increase in rank correlation from $91.6$ (direct evaluation) to $98.7$ (DISCO) in Table 3 translates into qualitative improvements in model ranking.
Figure 6 includes scatter plots of true vs. predicted ranks (see Figure 6). The direct-evaluation predictor demonstrates noticeable spread around the diagonal, while DISCO’s predictions align almost perfectly with it, indicating substantially more reliable ranking.
Figure 6: True vs. predicted rank comparison: direct evaluation vs. DISCO. $\rho_{s}$ means Spearman rank correlation.
Appendix D Additional Evaluation Results
D.1 Report Confidence Intervals
We report the standard deviation for the previously reported results on MMLU from Table 3, evaluated over one fixed chronological split and 5 independent runs.
We briefly remind the pipelines of the compared methods:
  • Selection: select a set of anchor points, i.e., a subset of the full test dataset based on different signals (random, IRT, model disagreement, etc.).
  • Prediction: estimate model performance on the full test dataset from outputs on anchor points.
That is why we use “Selection" and “Prediction" columns to explain the difference between methods. See § 4.1 for details on selection methods, and § 4.2 for details on prediction methods. MAE is mean absolute error; Rank is Spearman’s rank correlation.
Method Selection Prediction MAE $\downarrow$ Rank $\uparrow$
Baseline Random Direct evaluation $3.45\pm 0.67$ $91.6\pm 2.6$
tinyBenchmarks Random gp-IRT $2.79\pm 0.20$ $92.2\pm 2.3$
tinyBenchmarks Anchor-corr gp-IRT $2.08\pm 0.20$ $92.7\pm 2.1$
tinyBenchmarks Anchor-IRT gp-IRT $3.25\pm 0.49$ $92.2\pm 1.5$
DISCO High JSD KNN $1.14\pm 0.00$ $97.5\pm 0.0$
DISCO High JSD RF $1.30\pm 0.02$ $98.7\pm 0.1$
DISCO High PDS KNN $1.31\pm 0.00$ $97.2\pm 0.0$
DISCO High PDS RF $1.07\pm 0.04$ $98.7\pm 0.2$
Table 7: MMLU, single chronological split: MAE (%p) and Spearman rank (%), mean $\pm$ std over runs. DISCO (PDS/JSD + RF/kNN) has lower variance than baselines.
DISCO results are more stable than those of IRT and random sampling. This is because the only random component is Random Forest initialisation. In contrast, IRT is trained using variational inference, where stochastic gradient optimisation introduces additional randomness beyond model parameter initialisation.
D.2 Multiple Train/Test Split for Chronological Evaluation
To expand the number of chronological splits, we bootstrap 5 different train/test chronological splits using the following protocol: for each run, we split models into 385 old and 40 new based on timestamps, then bootstrap 346 source and 36 test models from these sets. Details on the chronological split can be seen in § 5.2. Results for the new splits can be seen below.
Method Selection Prediction MAE $\downarrow$ Rank $\uparrow$
Baseline Random Direct evaluation $2.85\pm 0.85$ $93.3\pm 3.0$
tinyBenchmarks Random gp-IRT $2.42\pm 0.43$ $93.6\pm 2.5$
tinyBenchmarks Anchor-corr gp-IRT $1.93\pm 0.31$ $92.9\pm 3.0$
tinyBenchmarks Anchor-IRT gp-IRT $3.13\pm 0.33$ $90.2\pm 4.5$
DISCO High PDS KNN $1.23\pm 0.09$ $97.0\pm 1.1$
DISCO High PDS RF $1.25\pm 0.14$ $98.0\pm 0.6$
Table 8: MMLU, bootstrapped chronological train/test splits (5 runs): MAE (%p) and Spearman rank correlation (%), mean $\pm$ std. DISCO remains best across methods.
Bootstrapped chronological splits slightly change the mean values (e.g., rank correlation from $98.6$ to $98.0$ and MAE from $1.06$ to $1.25$ for DISCO), but they do not alter the superiority of DISCO over other baselines.
Appendix E Sensitivity of DISCO to Model Calibration
To evaluate the sensitivity of DISCO to model calibration, we compared the Expected Calibration Error (ECE) of target models with the Mean Absolute Error (MAE) between their true performance and DISCO-predicted performance on MMLU. We observe a Pearson correlation of $0.49$ between MAE and ECE, indicating that better-calibrated models lead to more accurate performance (lower MAE) predictions by DISCO.
This phenomenon is explained by the information relationship between confidence and correctness. For a perfectly calibrated model, the mapping between prediction confidence and correctness is deterministic and monotonic, resulting in high mutual information. In contrast, for a highly miscalibrated model (e.g., random guessing or uniformly confident but incorrect), prediction confidence becomes statistically independent of correctness, leading to low mutual information. Consequently, the more calibrated a model is, the more predictive its confidence patterns are of its true performance, and therefore the more informative its signature is for DISCO performance prediction.
The corresponding scatter plot is shown in Figure 7.
Figure 7: Correlation between DISCO prediction error (MAE) and Expected Calibration Error (ECE).
During this analysis, we observed that two factors are confounded in calibration metrics: (1) overall confidence level, and (2) how well predictive uncertainty is reflected in confidence. To isolate the effect of overall confidence, we compared MAE with mean prediction confidence separately. We find a Pearson correlation of 0.47-0.47 between MAE and mean confidence, suggesting that overall model confidence is the dominant component of ECE that influences DISCO performance.
Figure 8 presents the corresponding scatter plot.
Figure 8: Correlation between DISCO prediction error (MAE) and mean model confidence.
Appendix F Performance Gap Experiments
In addition to source/target model splits discussed in § 5.4, we added experiments with a wider performance gap between source and target models to identify potential failure modes for DISCO. Inspired by (Zhang et al., 2025), we introduce a performance split with varying gaps. We sort all models by their average performance and take the top-10% or top-30% (40 or 128 models) as target models, while using the bottom-90% or bottom-50% (385 or 213 models) as source models. The accuracy gap between the weakest target model and the strongest source model is 0.07%p or 8.18%p.
All model splits are summarised in Table 9.
IID Chron. Performance w/o gap Perf. w/ gap
Prelim. model sorting By timestamp By performance By performance
Target models Every 10th model Top-10% Top-10% Top-30%
Source models Everything else Bottom-90% Bottom-90% Bottom-50%
Table 9: Source/target models splits.
Table 10 reports Spearman’s rank correlation.
IID Chron. Perf. w/o gap Perf. w/ gap
Direct eval on random subset $92.1\pm 1.3$ $91.6\pm 2.6$ $89.8\pm 5.9$ $87.4\pm 5.7$
DISCO $98.6\pm 0.3$ $98.7\pm 0.2$ $98.1\pm 0.2$ $89.2\pm 1.0$
Mean difference (DISCO – direct eval) +6.5 +7.1 +8.7 +1.8
Table 10: DISCO benefit vs direct evaluation on random subset across various source/target models splits.
For a source/target split with a performance gap, the difference between DISCO and direct evaluation is 1.8%p, which is lower than for the IID split (6.5%p), the chronological split (7.1%p), or the performance split without a gap (8.7%p).
While this scenario can be seen as a failure mode for DISCO, we believe that such a source/target performance gap does not happen in practice. Instead, the accuracies of source and target models are often mixed. There are two main reasons for that. First, when practitioners develop new models, their early versions are often worse than the best previously evaluated models. Second, it takes time before new models consistently outperform older ones. Infrequent extra evaluations can allow practitioners to always keep the performance gap low. That makes the performance split with a gap less realistic than other splits. We thus conclude that DISCO does not break when the source and target model distributions differ, but only when the difference is unrealistically substantial.
Appendix G Mutual Information and Jensen-Shannon Divergence